### v1

# Arboretum: A Large Multimodal Dataset Enabling AI for Biodiversity

Abstract
We introduce Arboretum, the largest publicly accessible dataset designed to advance AI for biodiversity applications. This dataset, curated from the iNaturalist community science platform and vetted by domain experts to ensure accuracy, includes 134.6 million images, surpassing existing datasets in scale by an order of magnitude. The dataset encompasses image-language paired data for a diverse set of species from birds (Aves), spiders/ticks/mites (Arachnida), insects (Insecta), plants (Plantae), fungus/mushrooms (Fungi), snails (Mollusca), and snakes/lizards (Reptilia), making it a valuable resource for multimodal vision-language AI models for biodiversity assessment and agriculture research. Each image is annotated with scientific names, taxonomic details, and common names, enhancing the robustness of AI model training.
We showcase the value of Arboretum by releasing a suite of CLIP models trained using a subset of 40 million captioned images. We introduce several new benchmarks for rigorous assessment, report accuracy for zero-shot learning, and evaluations across life stages, rare species, confounding species, and various levels of the taxonomic hierarchy.
We anticipate that Arboretum will spur the development of AI models that can enable a variety of digital tools ranging from pest control strategies, crop monitoring, and worldwide biodiversity assessment and environmental conservation. These advancements are critical for ensuring food security, preserving ecosystems, and mitigating the impacts of climate change. Arboretum is publicly available, easily accessible, and ready for immediate use. Please see the project website for links to our data, models, and code.
1 Introduction
AI advances are poised to play a crucial role in biodiversity conservation, ecology management, and agriculture. Already, AI tools have been shown to enable automated species identification, monitoring of ecological changes, and optimization of crop management [32, 5]. However, standard AI approaches for biodiversity applications persistently face major challenges. Training datasets are labor-intensive and costly to create; they cover only a narrow set of visual concepts; standard vision models excel at single tasks, but require extensive retraining for new tasks; models often struggle with generalizing to unseen labels and new environments, limiting their effectiveness in real-world applications [30, 13]. Models that perform well on benchmarks often fail in the wild [11, 1]. Standard computer vision datasets (ImageNet and its successors) have significant limitations, including incorrectly labeled images, geographical and cultural biases, and overlapping or ill-defined labels, all of which impair the development of high-performant AI models [20]. Consequently, there is a critical need for large, diverse, accurately annotated datasets that are specific to biodiversity, ecology, and agricultural research [23, 19].
In response to this need, several datasets have been introduced. Perhaps the most well-known (raw) pool of biodiversity images on the Web is iNaturalist [38], from which several curated datasets have been sourced, among them being iNat2021 [37] with 2.7M images of over 10,000 species of plants, animals, and fungi. However, insects (which comprise a very large fraction of extant species) are under-represented in this dataset. IP102 [40], Insecta [9], and the more recent BioScan-1M [12], are alternative datasets that focus on the Insecta Class. Perhaps the latest advance in such research is TreeOfLife-10M [35], which is currently the state-of-the-art dataset of text-annotated biological images, comprising 10M images with approximately 450K unique taxonomic classes.
Figure 1: The seven taxonomic classes in the Arboretum dataset. Example images of the top 5 most frequent species in each category, including their counts, common names, and scientific names. For clarity, the word ’class’ used in this context is denoted ’taxonomic class’ in the rest of the paper to distinguish it from the other common ML definitions.
In this paper, we make significant contributions to this body of work by curating and releasing Arboretum. This dataset includes 134.6 million captioned images of approximately 326.9K species. The dataset surpasses all existing datasets in scale by an order of magnitude, constituting the largest public, “AI-ready" dataset of curated biodiversity images. The dataset encompasses image-language paired data for a diverse set of species from birds (Aves), spiders/ticks/mites (Arachnida), insects (Insecta), plants (Plantae), fungus/mushrooms (Fungi), snails (Mollusca), and snakes/lizards (Reptilia). See Figure 1 for representative examples, and the project website for details.
Each image in Arboretum is sourced from the iNaturalist community science platform [38] and is annotated with the corresponding common (English) name, scientific (Latin) name, and taxonomic hierarchy. The metadata for each species has been vetted by domain experts, ensuring the accuracy of the text annotations for robust AI model training. We open-source the tooling pipeline used to curate Arboretum, and hope that the community will use this dataset as a valuable resource for further development of multimodal (vision-language) AI models for biodiversity and agriculture research.
To showcase the potential of Arboretum, we further make two technical contributions. First, we train and release ArborCLIP, a suite of vision-language foundation models trained on a subset of approximately 40M Arboretum samples, comprising approximately 33K species. This subset was constructed with standard filtering criteria in order to reduce skewness in species counts. Our models show excellent generalization performance and can support zero-shot (or few-shot) classification with either common or scientific names of unseen taxa. We also anticipate that these models can be fruitfully fine-tuned in the future on datasets for specific biodiversity-related applications.
Second, we rigorously quantify the performance of our foundation models on five existing fine-grained image classification benchmarks, as well on three newly curated test datasets. We find that ArborCLIP models comfortably achieve the state-of-the-art in certain settings, while both the original (OpenAI) CLIP model as well as BioCLIP [35] excel in certain other settings. We analyze these findings in further detail below, but overall we hope that our dataset can be used by the AI community as a testbed for further algorithmic and scaling research in fine-grained image recognition.
Figure 2: Distribution of the Arboretum dataset. (a) Size of each taxonomic class in the Arboretum dataset. (b) Species counts for each taxonomic class. (c) The 40 highest occuring species.
The remainder of this paper is organized as follows. Section 2 introduces the Arboretum dataset, the dataset’s salient characteristics, and a comparison with previous work. Section 3 details our curation methodology. Section 4 introduces our newly proposed test datasets and their characteristics. Section 5 details our new ArborCLIP models and their benchmark performance relative to previous work. Section 6 concludes with a discussion of limitations and potential future directions.
2 The Arboretum Dataset
Figure 3: Treemap diagram of the Arboretum dataset, starting from Kingdom. The nested boxes represent phyla, (taxonomic) classes, orders, and families. Box size represents the relative number of samples.
Characteristics.
Arboretum comprises over 134.6M images across seven taxonomic classes —Aves, Arachnida, Insecta, Plantae, Fungi, Mollusca, and Reptilia. These taxonomic classes were chosen to represent the span of species — outside of charismatic megafauna — which critically impact biodiversity. The images in Arboretum span 326,888 species. Overall, this dataset nearly matches the state-of-the-art curated dataset (TreeOfLife-10M) in terms of species diversity, while comfortably exceeding it in terms of scale by a factor of nearly 13.5×13.5$\times$13.5 ×. Figure 1 shows representative image samples, and Figure 2 displays the distribution of samples according to the seven categories (and the most frequently occurring species). Figure 3 displays phyla, taxonomic classes, orders, and families represented in the dataset.
Each image sample in Arboretum is annotated with a rich amount of curated metadata. This metadata design facilitates easy filtering of species by image count and taxonomic information. The metadata integrates common names, scientific names, and detailed taxonomic hierarchies. For the full list of fields, see Table 1.
Along with the dataset we also release our data curation tooling pipeline, that enables users to easily access and manipulate the dataset. The pipeline allows researchers to select specific categories, visualize data distributions, and manage class imbalance effectively based on their needs. It facilitates the downloading of specific images by their URLs and provides image-text pairs and user-defined chunks to support various AI applications. This pipeline enables users to define different subsets of Arboretum in an easy manner, overall making the dataset AI-ready and easy to use, and reducing the barriers to follow-up research in AI for biodiversity.
Dual-language text descriptions.
We adopt both common and scientific names, since Latin is a low-resource language and current vision-language backbones do not perform well on scientific names alone in a zero-shot manner. We found that a well-structured text description that integrate common names, scientific names, and detailed taxonomic hierarchies facilitates the learning of relationships between Latin and English terms, thereby improving the models’ applicability in scientific contexts [6, 34, 39]. Moreover, incorporating the taxonomic hierarchy enables models to more effectively associate visual data with taxonomic terminology [21, 2]. This matches the guidelines suggested by BioCLIP [35] to enhance model performance and generalization.
Table 1: Annotations provided in the Arboretum Dataset.
Text Type Description
Common Name Vernacular name (e.g., Western honey bee)
Scientific Name Genus and species (e.g., Apis mellifera)
Taxonomic Name Flattened taxonomy concatenated into a single string
Taxonomic Rank Specific level in the hierarchy (e.g., subspecies, species)
Privacy Measures: The images of Arboretum were sourced from the iNaturalist Open Dataset, whose metadata included Personally Identifiable Information (PII). This included information about observers, such as their usernames and sometimes their real names if they have chosen to share that information publicly. We removed all such fields to ensure that no PII is present in the metadata associated with Arboretum samples, ensuring the privacy of all contributors.
Offensive Content: Some of our URLs may point to images that users could find disturbing or harmful, such as photos of dead or dismembered animals. We retained these types of images since they sometimes can provide valuable scientific data about wildlife, including information on predation events, roadkill, and other occurrences relevant to conservation and biodiversity studies. Although iNaturalist relies on user contributions and community moderation to maintain the quality and appropriateness of the data, we acknowledge that the vast and diverse nature of the data means that some offensive or inappropriate content might be present.
Our closest comparisons are with BioScan-1M (which appeared in NeurIPS 2023 Datasets and Benchmarks) and TreeOfLife-10M (which will appear in CVPR 2024). BioScan-1M focuses solely on the Insecta Class and provides scientific names, taxonomic ranks, as well as DNA barcodes. The TreeOfLife-10M dataset comprises 10.4 million images, integrating data from iNat2021 [37], BioScan-1M, and a fresh set of image samples sourced from the Encyclopedia of Life (EOL). It also supports dual-language labels and detailed taxonomic hierarchies, and was used to train the BioCLIP vision-language model. See Table 2 for essential differences.
Table 2: Comparison of Arboretum with existing biodiversity datasets.
Feature Arboretum TreeOfLife BioScan
Size 134.6 million images 10.4 million images 1.1 million images
Diversity 326.9K species 454.1K species 8.3K
Labels Provided Dual language (common and scientific names), detailed taxonomic hierarchies Dual language (common and scientific names), detailed taxonomic hierarchies Single language (scientific names), taxonomic ranks (family to species), DNA barcodes
Data Source iNaturalist Open Dataset iNaturalist, Encyclopedia of Life (EOL), BioScan-1M Specimens from Malaise traps, DNA barcodes matched to BOLD
Key Features Ready-to-use pipeline, reduce class imbalance, high-quality annotations, supports ArborCLIP Rich hierarchical representations, comprehensive metadata, supports BioCLIP Focus on insects, high-resolution images, detailed taxonomic annotation, DNA codes
3 Data Collection and Curation Methodology
Challenges with iNaturalist Open Data.
All of Arboretum is sourced from the iNaturalist Open Data community science platform, which (in all) comprises over 190M biodiversity-relevant observations shared by users. However, there are still significant gaps in usability for AI research. The photos and metadata, although easily downloadable, are provided in four separate metadata sheets that are not ready to use. Taxa information is encoded as numerical IDs, requiring additional API calls and non-trivial lookups to convert these into common or scientific names. The multiple metadata sheets structure is fragmented across four separate files—photos, taxa, observations, and observers—adding complexity to data integration. Managing data balance and filtering out species with too few images can lead to biases towards common (charismatic) species and an imbalanced training process.
Curation of Arboretum.
The iNaturalist Open Dataset comprises a collection of 250.0M images stored on an AWS S3 bucket as of May 27, 2024, with associated metadata in the form of four separate CSV files (indexed as photos, observations, taxa, and observers). Details of each of these files are provided in Section A.5 in the Appendix. While these files contain a plethora of interesitng information, they are designed for rapid lookup, and therefore not “AI-ready". To resolve this, we curate the metadata into a streamlined format for easy usage by AI practitioners.
We first use the corresponding CSV files to populate an SQL database with each CSV file as its own SQL table. We create a new aggregate SQL table by joining photos, observations, and taxa on its relational columns, and discarding columns not relevant to us. We also construct a new column in the aggregated table and populate it with the Amazon S3 URL where the image file is hosted.
We then query this SQL table and extract rows corresponding to the following seven Arboretum categories: Aves, Arachnida, Insecta, Plantae, Fungi, Mollusca, and Reptilia. [^1] As mentioned above, we choose these 7 classes since they represent non-megafauna that criticially impact biodiversity. Typically, megafauna species are well suited to standard image recognition models, but there remains a pressing need for similar efforts for these particular classes. The iNaturalist Open Data metadata files do not contain common name information; to reconstruct this information, we cross-match species names from the iNaturalist Taxonomy DarwinCore Archive, which is compiled monthly. Once the common name is reconstructed we append this to the SQL table. Finally, we export this table as a CSV file, which we release for public use on HuggingFace.
Data filtering and preprocessing.
As described above, Arboretum comprises well-processed metadata with full taxa information and URLs pointing to image files. The metadata can be used to filter specific categories, visualize data distribution, and manage imbalance effectively. We provide a collection of software tools that enable users to easily download, access, and manipulate the dataset. This makes it straightforward for researchers to explore different data subsets.
The citizen science aspect of iNaturalist data leads to a wide variation in the number of observations per species, with some species having only a few records while others have thousands. To mitigate this, our tools can be used to apply user-defined filtering criteria to exclude species with fewer than a specified number of images. Additionally, we imposed an upper limit on the number of images per species to mitigate overfitting.
To address imbalances during training (described further below in our experiments), we download the data and organize into chunked tar files using a semi-global shuffling strategy. Initially, each tar file is shuffled and divided into smaller groups. These groups were then randomly merged to form larger batches, ensuring a balanced distribution of species within each batch. This approach significantly improves dataset integrity by preventing any single species from spanning full batches.
4 Models and Benchmarks
We now showcase demonstrate the utility of the Arboretum dataset by creating and benchmarking ArborCLIP, a new suite of vision-language foundation models for biodiversity.
4.1 Arboretum-40M
To start, we first construct Arboretum-40M, a subset comprising approximately 40M samples and 33K species, which we constructed by applying the data filtering tools described above.
The data included in this subset include all samples in the seven Arboretum categories posted on iNaturalist prior to January 27, 2024. We applied filtering criteria to exclude species with fewer than 30 images and imposed a maximum limit of 50,000 samples per species. We conducted semi-global shuffling and divided the data into mini-batches of approximately 50,000 samples each. From these mini-batches, 95% were randomly selected for training and validation, while the remaining 5% were reserved for testing. Detailed information can be found in Table 3.
Table 3: Training data sources used in Arboretum-40M and Diversity in Different Taxonomy Levels. We integrate taxonomic labels into the images.
Dataset Description Images Unique Classes TreeOfLife-10M Dataset combines a subset of iNaturalist, Encyclopedia of Life (EOL), BioScan-1M. 10.4M 454,103 Arboretum-40M One subset of Arboretum with size 40M. 39.9M 33,364 Level Uniques kingdom 3 phylum 14 class 50 order 311 family 1692 genus 11506 species 33364
4.2 New Benchmarks
From Arboretum, we create three new benchmark datasets for fine-grained image classification. In addition, we report results on several established benchmarks in the literature; see Table 4.
Arboretum-Balanced. To enforce a balanced species distribution across the seven categories, we curate Arboretum-Balanced. Each category includes up to 500 species, with 50 images per species. Table 6 in the Appendix shows each category’s exact number of species.
Arboretum-Unseen. To provide a robust benchmark for evaluating the generalization capability of models on unseen species, we curate Arboretum-Unseen. The test dataset was constructed by identifying species with fewer than 30 instances in Arboretum, ensuring that the dataset contains species that were unseen by ArborCLIP. Each species contained 10 images.
Arboretum-LifeStages. To assess the model’s ability to recognize species across various developmental stages, we curate Arboretum-LifeStages (see Figure 4). This dataset has 20 labels in total and focuses on insects, since these species often exhibits significant visual differences across their lifespan. Arboretum-LifeStages contains five insect species and utilized the observation export feature on the iNaturalist platform to collect data from 2/1/2024 to 5/20/2024 to ensure no overlap with the training dataset. For each species, life stage filters (egg, larva, pupa, or adult) were applied.
Figure 4: (a) Example images from Arboretum-Unseen. (b) Arboretum-Life-Stages with 20 class labels: four life stages (egg, larva, pupa, and adult) for five distinct insect species.
4.3 ArborCLIP: New vision-language foundation models for biodiversity
We use Arboretum-40M to train new CLIP-style foundation models, and then evaluate them on zero-shot image classification tasks. Following the implementation of Stevens et al. [35], we utilize a ViT-B/16 architecture initialized from the OpenAI CLIP weights [29], and train for 40 epochs. We also train a ViT-L/14 model from the MetaCLIP [41] checkpoint for 12 epochs, and a ViT-B/16 from the BioCLIP checkpoint for 8 epochs. All training hyperparameters are included in the Appendix (Section B). We compare with OpenAI’s ViT-B/16 CLIP model, the BioCLIP ViT-B/16 checkpoint, and MetaCLIP-CC ViT-L/14. We publicly release all code needed to reproduce our results here.
5 Experimental Results
Table 4: Existing benchmark datasets; our novel datasets are described separately in section 4.2.
Name Description Examples Classes Labels
Anim Birds 525 Scraped dataset of bird images from web search. [27] 89 8858988589\,88589 885 $525$ Taxonomic
BioCLIP-Rare Subset of species in the IUCN Red List categories: Near Threatened through Extinct in the Wild (iucnredlist.org). 12 0001200012\,00012 000 $400$ Taxonomic
Plt & Fun Fungi Expert-labeled images of Danish fungi [26]. $1000$ $25$ Scientific
DeepWeeds Weed images collected in situ from eight rangelands across northern Australia [25] 17 5091750917\,50917 509 $9$ Common
Inse Confounding Species Dataset evaluating models on challenging visually similar species pairs [4]. $100$ $10$ Mixed
Insects-2 Mixed common and scientific name classification for insect pests [40]. $4080$ $102$ Mixed
Metrics. For each of the benchmark datasets, we report top-1 zero-shot accuracy. For datasets where taxonomic information is available, we report accuracy over scientific names. For datasets where no taxonomic information is available, we utilize category names provided by the benchmark’s authors. For our Arboretum-Life-Stages benchmark, we consider a class label to be a unique (species | life-stage) tuple, resulting in a 20-class benchmark. Finally, we report an aggregate metric which is the weighted average over unique class labels across all benchmarks in our suite.
Overview of results. In Table 5, we report the results of our core benchmark suite. At a high level, we observe that ArborCLIP variants achieve the best accuracy averaged over benchmarks. In particular, they perform extremely well on Arboretum-Balanced (a remarkable 91.1 top-1 accuracy over 2250+ class labels). ArborCLIP also does very well on the Fungi dataset (even though the Fungi class is not central to Arboretum-40M), and the DeepWeeds dataset. Therefore, ArborCLIP exhibits strong generalization capabilities across diverse datasets.
We also observe that BioCLIP performs very well on Arboretum-Unseen and BioCLIP-Rare. The reasons might be that BioCLIP has seen approximately 450K species, and there might be nontrivial overlap with the species set in Arboretum-Unseen. On the other hand, it could be that ArborCLIP suffers from forgetting issues while training on Arboretum-40M. For BioCLIP-Rare, the dataset is a subset from EOL which BioCLIP did not see before, but TreeofLife contains the majority of the EOL dataset.
Limitations.
We also evaluated all models on the challenging Confounding-species benchmark introduced in [4], but find that all models perform at or below random chance, and do not report results here; this could be an interesting avenue for follow-up work.
In Table 7 in the Appendix, we report model performance at different levels of the taxonomic hierarchy. Generally, we find that models trained on web-scraped data perform better with common names, whereas models trained on specialist datasets perform better when using scientific names. Additionally, models trained on web-scraped data excel at classifying at the highest taxonomic level (kingdom), while models begin to benefit from specialist datasets like Arboretum-40M and Tree-of-Life-10M at the lower taxonomic levels (order and species). From a practical standpoint, this is not problematic: ArborCLIP is highly accurate at the species level, and higher-level taxa can be deterministically derived from lower ones.
Table 5: ArborCLIP performs well on a range of benchmarks. ArborCLIP-O was pretrained from an OpenAI model checkpoint, ArborCLIP-B from the BioCLIP checkpoint, and ArborCLIP-M from a MetaCLIP-cc checkpoint. AU stands for Arboretum-Unseen, using Scientific Names (n-classes=300). AB stands for Arboretum-Balanced, using Scientific Names (n-classes=2253). BCR stands for the BioCLIP Rare species benchmark (n-classes=400). F stands for Fungi (n-classes=25). I2 stands for Insects2 (n-classes=102). B stands for Birds525 (n-classes=525), LS for the Life Stages benchmark (n-classes=20), and DW for DeepWeeds (n-classes=9).
Dataset Arch AU AB BCR F I2 B LS DW Weighted Average
OpenAI ViT-B-16 12.9 7.3 10.9 11.5 10.2 50 56.5 10.3 14.7
BioCLIP ViT-B-16 68.2 62.2 30.2 45.1 20.8 68.7 18 19.9 58.5
MetaCLIP-cc ViT-L-14 24.9 15.4 20.5 24.6 16.1 70.1 64.3 14.7 25.0
ArborCLIP-O ViT-B-16 47.1 91.1 22.9 43.2 16.5 47.8 28 17 70.8
ArborCLIP-B ViT-B-16 53.8 82.2 23.7 53.9 16.9 57.1 15 18.4 67.2
ArborCLIP-M ViT-L-14 44.3 91.1 21.8 54.7 5.1 42.5 26.3 49.9 69.5
Addressing these limitations will further enhance the applicability of models like ArborCLIP in real-world biodiversity monitoring tasks.
6 Concluding Discussion
We introduce Arboretum, the largest publicly accessible dataset designed to advance AI for biodiversity applications. This dataset, curated from the iNaturalist community science platform, includes 134.6 million images, surpassing existing datasets in scale by an order of magnitude. We anticipate that Arboretum will enable the development of AI models that can enable various digital tools ranging from pest control strategies, crop monitoring, and worldwide biodiversity assessment and environmental conservation.
We also believe that Arboretum can be used as a unique testbed for measuring progress on fine-grained image recognition. The success of ArborCLIP on Arboretum-Unseen underscores the importance of scaling up per-category sample size, or vertical scaling [8], in achieving high accuracy on long-tailed extreme-imbalance classification. However, BioCLIP continues to exhibit superior performance on several datasets, and we believe that this is because TreeofLife-10M contains an order-of-magnitude more classes (species) than Arboretum-40M. We invite the AI community to create new subsets of Arboretum with varying degrees of balance and species diversity, and use our tooling to measure model performance against current benchmarks.
Acknowledgements
This work was supported by the AI Research Institutes program supported by the NSF and USDA-NIFA under AI Institute: for Resilient Agriculture, Award No. 2021-67021-35329. This was also partly supported by the NSF under CPS Frontier grant CNS-1954556. CH, BF, AN, and KM gratefully acknowledge the support of NYU IT High Performance Computing resources, services, and staff expertise.
Appendix A Appendix
A.1 Background on CLIP and zero-shot classification
Unlike traditional vision models, CLIP jointly trains an image encoder and a text encoder to predict the correct pairings of a batch of (image, text) examples, leveraging natural language supervision to enhance generalization [29]. CLIP’s approach allows it to learn from a wide variety of images and their associated textual descriptions, making it more flexible and general compared to standard vision models. This flexibility is crucial for in various domains, including biodiversity monitoring and agriculture. For instance, CLIP models analyze digital plant specimen images, aiding in pre-processing and filtering for further analysis for agriculture purposes [15, 16]. As for biodiversity, WildCLIP and KI-CLIP facilitate wildlife observation and monitoring with high accuracy and effectiveness in data-sparse settings [10, 22]. These examples underscore the importance of developing and utilizing comprehensive datasets to fully leverage the capabilities of CLIP models in advancing biodiversity and agricultural research.
A.2 The value of taxonomic information
Taxonomic classification, the hierarchical arrangement of organisms into categories based on shared characteristics, is foundational in biological sciences. Taxonomy underpins various scientific, ecological, and agricultural applications. It allows for precise identification and classification of species, which is fundamental for understanding biodiversity and monitoring ecosystems. For instance, accurate species identification can aid in tracking invasive species, as noted in studies such as [33]. In agriculture, detailed taxonomic information helps in identifying pests and beneficial species, thereby improving pest control strategies and crop management; supports ecological research by providing insights into species interactions, distribution patterns, and evolutionary relationships [12]; and is essential for policy-making and conservation planning [31].
A.3 Scientific versus common names
Although we identify the importance and need to include taxonomic information in the dataset for biodiversity, one potential challenge is the fact that this information is mostly in Latin for which text embedding models often exhibit suboptimal performance due to its status as a low-resource language [36]. Nonetheless, Latin remains indispensable as it is the standard for representing scientific names and taxonomic classifications. We therefore integrate common names, scientific names, and detailed taxonomic hierarchies. We believe that such an “all-encompassing” approach facilitates the learning of relationships between Latin and English terms, thereby improving the models’ applicability in scientific contexts [6, 34, 39]. Furthermore, incorporating taxonomic data into the training process significantly enhances the multimodal capabilities of the models, enabling them to associate visual data with taxonomic terminology [21, 2].
A.4 iNaturalist, iNaturalist Open Data
iNaturalist is an online social network for sharing biodiversity information and learning about nature. It serves as a crowdsourced species identification system and organism occurrence recording tool. Users from around the world upload images, making the continuously updated dataset valuable for AI applications in biodiversity and research. Each photo includes detailed metadata: copyright status, location, uploader, time, and taxonomic classification. This diversity in image sources makes iNaturalist an excellent dataset for training AI models intended for real-world applications [37, 24, 7, 3]. Despite its vast and diverse data, iNaturalist is not directly optimized for AI researchers: arranging this data for use in AI models like CLIP is not straightforward. Each photo has its own page on the iNaturalist website, making it difficult to download images along with all the necessary information in a streamlined manner.
The iNaturalist Open Dataset aims to address some of these challenges. It is one of the world’s largest public datasets of photos of living organisms, structured as a "bucket" of images stored using Amazon Web Service’s Simple Storage Service (S3). The dataset includes multiple resized versions of each photo, allowing users to download the size most useful to their research.
Additionally, the dataset provides four tab-separated CSV files representing observations, observers, photos, and taxa_id. These files are generated monthly, capturing a snapshot of the continually changing iNaturalist data. The images in the iNaturalist Open Dataset are licensed under either CC0, CC-BY, or CC-BY-NC and are open for public research. Photos with a CC0 license can be attributed as "[observer name or login], no rights reserved (CC0)". Photos with other Creative Commons licenses can be attributed as "© [observer name or login], some rights reserved ([license abbreviation])".
A.5 iNaturalist Details
Each image in the iNaturalist Open Dataset can be associated with its appropriate metadata through a group of four metadata CSV files,
representing photos, observations, taxa, and observers.
The photos metadata file contain nine distinct columns of metadata information of each photo. Of these columns, only photo_id and observation_uuid are relevant for us. The value of photo_id is a identifier number used to access individual photos, the photo’s iNaturalist page can be found by constructing a URL in this format: https://www.inaturalist.org/photos/[photo_id]. The value of observation_uuid indicates which observation the photo is associated with, it is used to map the photos metadata to the observations metadata.
An observation represents one user submission of a species encounter to the iNaturalist website. One observation can have multiple photos of the same species but never multiple species. The observation metedata file contains eight distinct columns of metadata information on each observation. The columns relevant to us are observation_uuid, quality grade, and taxon_id. Each observation is given a unique number identifier indicated by its observation_uuid. iNaturalist has its own system to determining the quality of an observation and its associated photos, quality_grade represents this and can range from "Casual", "Research Grade", or "Needs ID". The value taxon_id indicates the species is represented in the observation, it is used to map the observations metadata to the taxa metadata.
[…]
The observer metadata file comtains information about each user within the iNaturalist site. For the purpose of machine learning research none of its three metadata columns are relevant.
While the iNaturalist Open Dataset metadata files provide a plethora of interesting information, its structure makes it inherently cumbersome to use for research. To solve this, we aggregate and process the iNaturalist metadata into a concise and streamlined format for easy query and usage.
First, the respective CSV files are used to populate a SQL database with each CSV file as its own SQL table. A new aggregate SQL table is created that joins the photos, observations, and taxa tables on its relational columns. Only the metadata columns we deemed relevant are kept and the extraneous non-useful metadata columns are discarded.
One of the difficulties working with the base iNaturalist metadata files is that it does not contain the image URL, information that is critical in image downloads. We include a new column in the aggregated metadata table that explicitly links to the Amazon S3 URL in which the image is hosted.
The Arboretum metadata file used for model training contains the metadata columns phylum, class, order, family, genus, species, scientific_name, common_name for the seven Arboretum categories Aves, Arachnida, Insecta, Plantae, Fungi, Mollusca, and Reptilia. To ensure that only images and metadata from the seven Arboretum categories appear in our final dataset we use the taxa table to find the taxon in our categories then use it in a SQL query on the ancestry column of our aggregated metadata table.
The taxonomic rank columns are also found utilizing the ancestry metadata column. A difficulty in working with the ancestry metadata is present in that there is not a clear indication of what taxonomic rank a taxon id represents the ancestry string. This problem is exacerbated due to the presence of taxonomic ranks and dsub ranks whose presence is variable across different species. As such, a custom function is applied to each row to dynamically find the rank of each taxon id in the ancestry and then appropriately populate the taxon id to a metadata column of that rank. This process results in all taxonomies rank represented as metadata columns; only phyllum, class, order, family, genus and species are kept in the Arboretum metadata file.
The scientific name of a species is found using the name metadata column of our aggregated metadata table. The common name of a species is also useful metadata information. Unfortunately, the iNaturalist Open Data metadata files do not contain the common name information of a species. To address this, we curate a lookup table of the common names in our dataset. This is obtained from the iNaturalist Taxonomy DarwinCore Archive, Having obtained the common names for each species, we append it to the Arboretum-specific metadata.
A.6 Composition of Arboretum-40M
See Figure 5 and Table 6.
Figure 5: Arboretum-40M Dataset Analysis: a) Consistent category distribution across Arboretum-40M and Arboretum-116M datasets. b) Species exhibit a long-tailed distribution. c) Impact of local vs. semi-global shuffling on species representation within training minibatches.
Table 6: Number of Unique Species in Each Category in Arboretum-Balanced.
Category Number of Unique Species
Kingdom: Fungi 281
Kingdom: Plantae 500
Phylum: Mollusca 147
Class: Insecta 500
Class: Arachnida 136
Class: Reptilia 189
Class: Aves 500
Appendix B ArborCLIP training details
We use Arboretum-40M to train new CLIP-style foundation models, and then evaluate them on zero-shot image classification tasks. Following the implementation of Stevens et al. [35], we utilize a ViT-B/16 architecture initialized from the OpenAI pretrained weights for our main model, and train for 40 epochs.
In addition, we also train a ViT-L/14 model from the MetaCLIP [41] checkpoint for 12 epochs, and a ViT-B/16 from the BioCLIP checkpoint for 8 epochs. We select the AdamW optimizer from Loshchilov and Hutter [18] along with a cosine learning rate scheduler, as this has previously been shown to perform well for CLIP pretraining [28]. We conduct twenty rounds of hyperparameter optimization using Ray Tune [17] to determine the optimal learning rate, $\beta_{1}$, $\beta_{2}$ and weight decay settings.
We train our models for a combined 10 days on 8xH100 nodes in bfloat16 precision [14] with gradient checkpointing, computing loss with local features, and utilizing static graph optimization for DDP.
Appendix C Additional ArborCLIP results
In Table 7, we report model performance at different levels of the taxonomic hierarchy. Generally, we find that models trained on web-scraped data perform better with common names, whereas models trained on specialist datasets perform better when using scientific names. Additionally, models trained on web-scraped data excel at classifying at the highest taxonomic level (kingdom), while models begin to benefit from specialist datasets like Arboretum-40M and Tree-of-Life-10M at the lower taxonomic levels (order and species).
However, ArborCLIP shows a performance decline at taxonomic levels below the species level. This is likely because our training metadata structure allows for classifications solely by referring to species information. From a practical standpoint, this is not problematic for the species in our test set since ArborCLIP is highly accurate at the species level, and higher-level taxa can be deterministically derived from the lower ones.
Furthermore, the OpenCLIP and MetaCLIP baselines outperform ArborCLIP on the life stages benchmark. This highlights the importance of retaining the general linguistic capabilities of the pretrained CLIP models for hybrid tasks.
Table 7: Different levels of taxonomic classification reveal the strengths and weaknesses of different models. DS stands for dataset, Arch for model architecture AU stands for Arboretum-Unseen, AB stands for Arboretum-Balanced, BCR for BioCLIP-Rare, K for Kingdom taxonomy, O for Order taxonomy, and CN for species taxonomy using common names.
DS Arch AU-K AU-O AU-CN AB-K AB-O AB-CN BCR-CN
OpenAI ViT-B-16 78 26.7 18.6 79.9 26.7 13.1 31.6
BioCLIP ViT-B-16 75.5 38.4 41.8 73.2 34.5 42.7 32.1
MetaCLIP-cc ViT-L-14 90.3 33.2 27.9 87.1 34.5 19.9 42.4
ArborCLIP from OpenAI ViT-B-16 62.9 13.5 29.7 56.5 12.4 70.8 22.2
ArborCLIP from BioCLIP ViT-B-16 62.8 17.8 24 65.7 14.5 52.6 23.6
ArborCLIP from MetaCLIP ViT-L-14 68.3 9.6 4.3 58.1 3.8 15.7 8.8